# Beyond the Destination: A Novel Benchmark for Exploration-Aware Embodied Question Answering

### v2

@@ 3.2 Dataset Generation Pipeline
Stage 3: Data Filtering. Although large models perform well, their outputs are not always reliable and reasonable. Additionally, due to the vast environment, there may be multiple rooms of the same type (e.g., bedrooms) or duplicate objects in the scene, which can result in non-unique answers to the questions. Therefore, manual screening of the generated data is necessary. First, we ensure that the answers to all questions remain relevant to the scene. Second, we utilize the generated trajectory video and the top-down views of each floor provided by the Habitat simulator to track the agent’s position and movement in the scene. This allows us to address issues where the target region cannot be clearly identified or reached accurately during exploration. Specifically, we retain only those questions where either no other regions of the same type as the target region, or where the target region is the closest to the initial position among similar regions, ensuring the uniqueness of the answer. When necessary, we also specify the key regions of the questions to guarantee that each expected answer is unique. We invited several individuals to review the data. To ensure the dataset’s high quality, a curator performed a final check to verify its overall consistency.
@@ 5.1 Baselines
Socratic Models. Using ground truth data $G$, we simulate an agent’s navigation and exploration to obtain first-person visual observations, from which we randomly sample $k$ frames. Each frame $G_{i}$ is described by an image captioning model as $d_{i}=\textrm{Captioner}(Q,G_{i})$, which are then concatenated into $D=[d_{1},d_{2},...,d_{k}]$. We set $k=\lceil\frac{L}{10}\rceil$, where $L$ denotes the step length of the ground truth. The LLMs then generate answers based on the question and the concatenated descriptions, i.e., $A=\textrm{LLMs}(Q,D)$. For caption generation, we employ LLaVA-v1.5-7b, and for the LLMs, we use DeepSeek-V3, GPT-4 and LLaMA-3-8b.
Multi-Frame VLMs. We randomly selecting k frames from the ground truth, which are subsequently fed into the VLMs, i.e., $A=\textrm{VLMs}(Q,G_{1},G_{2},...,G_{k})$. We employ GPT-4o-mini[15] and LLaVA-v1.5-7b.
@@
(4) Fine-EQA (Ours). For LLMs, we use GPT-4 to identify task-related functional regions. For VLMs, we employ Prismatic VLMs [19] to assign semantic values and evaluate functional regions. GPT-4o-mini is used to decide when to stop exploration and generate responses.
@@ 5.2 Comparison with State-of-the-art Methods
Non-Exploratory Agents. Among LLMs without visual information, DeepSeek-V3 demonstrates the best performance. When combined with Llava-1.5-7B for reasoning, their performance is improved, with DeepSeek-V3-based Socratic agent yielding the highest results. However, since EXPRESS-Bench focuses on active exploration, with questions typically related to the last frame, the performance improvement of the Socratic Models, which relies solely on passively observed video frames, remains limited. Additionally, Multi-Frame VLMs performs worse than the Socratic models, indicating that critical information extraction is essential for exploration-aware EQA than directly using images. This supports the reliability of EXPRESS-Bench in evaluating active exploration capabilities.
@@
Table 3: Ablation results.
@@
5.3 Ablation Studies
@@ 8.2.1 Formulas for Metric Calculation
where $ce_{i}$ represents the confidence of the VLM for the image, $l_{i}$ represents the distance the agent navigate along the ground truth path that is sufficient to complete the task, and $p_{i}$ is the actual distance the agent moves during the experiment.
@@ 8.2.2 Comparison of Metrics
Figure 8: Comparison of the metrics proposed by OpenEQA and ours. By considering the grounding of the response, our metric offers a more reliable assessment of the model’s performance.

### v3

@@ 3.2 Dataset Generation Pipeline
Stage 3: Data Filtering. Although large models perform well, their outputs are not always reliable and reasonable. Additionally, due to the vast environment, there may be multiple rooms of the same type (e.g., bedrooms) or duplicate objects in the scene, which can result in non-unique answers to the questions. Therefore, manual screening of the generated data is necessary. First, we ensure that the answers to all questions remain relevant to the scene. Second, we utilize the generated trajectory video and the top-down views of each floor provided by the Habitat simulator to track the agent’s position and movement in the scene. This allows us to address issues where the target region cannot be clearly identified or reached accurately during exploration. Specifically, we retain only those questions where either no other regions of the same type as the target region, or where the target region is the closest to the initial position among similar regions, ensuring the uniqueness of the answer. When necessary, we also specify key regions of the questions and include contextual details, such as the location, attributes, and relationships of the target object. We invited several individuals to review the data. To ensure the dataset’s high quality, a curator performed a final check to verify its overall consistency.
@@ 5.1 Baselines
Socratic Models. Using ground truth data $G$, we simulate an agent’s navigation and exploration to obtain first-person visual observations, from which we randomly sample $k$ frames. Each frame $G_{i}$ is described by an image captioning model as $d_{i}=\textrm{Captioner}(Q,G_{i})$, which are then concatenated into $D=[d_{1},d_{2},...,d_{k}]$. We set $k=\lceil\frac{L}{10}\rceil$, where $L$ denotes the step length of the ground truth. The LLMs then generate answers based on the question and the concatenated descriptions, i.e., $A=\textrm{LLMs}(Q,D)$. For caption generation, we employ GPT-4o-mini[15] and LLaVA-v1.5-7b, and for the LLMs, we use DeepSeek-V3, GPT-4 and LLaMA-3-8b.
Multi-Frame VLMs. We randomly select k frames from the ground truth, which are subsequently fed into the VLMs, i.e., $A=\textrm{VLMs}(Q,G_{1},G_{2},...,G_{k})$. We employ GPT-4o-mini and LLaVA-v1.5-7b.
@@
(4) Fine-EQA (Ours). For LLMs, we use GPT-4 to identify task-related functional regions. For VLMs, we employ Prismatic VLMs (Prism-DINOSigLIP 7B) [19] to assign semantic values and evaluate functional regions. GPT-4o-mini is used to decide when to stop exploration and generate responses.
@@ 5.2 Comparison with State-of-the-art Methods
Non-Exploratory Agents. Among LLMs without visual information, DeepSeek-V3 demonstrates the best performance. When combined with VLMs for reasoning, their performance is improved, with DeepSeek-V3-based Socratic agents yielding the highest results. However, since EXPRESS-Bench focuses on active exploration, with questions typically related to the last frame, the performance improvement of the Socratic Models, which relies solely on passively observed video frames, remains limited. Additionally, Multi-Frame VLMs performs worse than the Socratic models, indicating that critical information extraction is essential for exploration-aware EQA than simply using numerous images. This supports the reliability of EXPRESS-Bench in evaluating active exploration capabilities.
@@
5.3 Experiments on Other Datasets
To further evaluate the performance of Fine-EQA, we conduct experiments on two additional datasets. For the OpenEQA[29], we focus specifically on the A-EQA subset, which assesses the agent’s ability to explore the environment and answer questions. We use the corresponding evaluation metrics $C^{\prime}$ and $E^{\prime}$ for performance measurement, with their formulas provided in the Supplementary Material.
The results are presented in Tab.3. Fine-EQA outperforms the best-performing GPT-4V model from [29], particularly in terms of exploration efficiency. This is because the active exploration strategy in [29] relies entirely on the frontier-based method and fails to terminate exploration promptly after gathering the information necessary for the task.
Table 3: Performance comparison on the A-EQA subset of OpenEQA. Results marked with * are from the OpenEQA benchmark, where GPT-4V is evaluated on a random subset of 184 questions. In contrast, our Fine-EQA is evaluated on the full set of questions.
$C^{\prime}$ $E’$ ’
OpenEQA w/ GPT-4V 41.8±3.2* 7.5±0.6*
Fine-EQA 43.27 29.16
We also evaluated the performance of Explore-EQA[31] and Fine-EQA on the multiple-choice dataset HM-EQA[31], using answer accuracy and the length of the agent’s navigation path as metrics. As shown in Tab.4, Fine-EQA achieves substantial improvements over Explore-EQA in both metrics.
Table 4: Performance comparison on HM-EQA.
Accuracy(%)↑ Path Length(m)↓
Explore-EQA 50.4 93.687
Fine-EQA 56.0 54.267
Furthermore, we analyzed the ambiguity rate in questions, finding 35.6% in HM-EQA, 62.3% in Open-EQA, and only 9.7% in EXPRESS-Bench. The absence of complete exploration trajectories or clear target locations in HM-EQA and Open-EQA further hinders accurate performance evaluation. In contrast, our EXPRESS-Bench enhances evaluation reliability and interpretability.
5.4 Ablation Studies
Table 5: Ablation Study of Model Modules.
@@ 5.4 Ablation Studies
To isolate the impact of GPT-4o-mini on Fine-EQA’s performance, we replaced it with other VLMs and conducted evaluations on the EXPRESS-Bench. As shown in Tab.6, Fine-EQA built using different VLMs exhibit varying performance across different metrics, but consistently outperform other models. We observed that Fine-EQA models with higher $C$ scores tend to engage in more extensive exploration within the environment, which is reflected in their lower $E$ scores.
Table 6: Ablation Study of VLMs.
$C$ $C^{*}$ $E_{path}$ $d_{T}$
Janus-Pro-7B 39.88 63.35 20.86 6.18
Qwen-Vl-Plus 40.02 63.06 17.57 6.41
GPT-4o-mini 40.55 63.95 16.22 6.43
@@ 8.2.1 Formulas for Metric Calculation
The calculation formulas for metrics on the OpenEQA are as follows:
$C^{\prime}=\frac{1}{N}\sum\limits_{i=1}^{N}\frac{{\sigma}_{i}^{{}^{\prime}}-1}{4}\times 100\%$ (12)
$E^{\prime}=\frac{1}{N}\sum\limits_{i=1}^{N}\frac{{\sigma}^{{}^{\prime}}_{i}-1}{4}\times\frac{l_{i}}{max(p_{i},l_{i})}\times 100\%$ (13)
where $\sigma_{i}^{{}^{\prime}}$ is determined by the LLMs based on the OpenEQA prompt, $l_{i}$ represents the distance the agent navigate along the ground truth path that is sufficient to complete the task, and $p_{i}$ is the actual distance the agent moves during the experiment.
@@
where $ce_{i}$ represents the confidence of the VLM for the image.
Figure 8: Comparison of the metrics proposed by OpenEQA and ours. The EAC metric combines $\sigma$ and $\delta$ to jointly assess both the semantic validity and visual grounding of the response. By considering the grounding of the response, our metric offers a more reliable assessment of the model’s performance.